3 moles of hydrogen occupy a volume of 67.20 L, at standard temperature and pressure (0 °C and 760 mmHg). what is the density of hydrogen at STP?
Answer #1

moles of hydrogen=3 moles

mass of hydrogen= moles*molar mass

=3*2= 6 grams

Density=Mass/ volume

= 6/67.2

=0.089 g/L

Please like the solution.Feel free to ask any doubts in comments.